# Dirac Fermions and Topological Phases in Magnetic Topological Insulator Films

### v2

@@ VI.1.3 Stronger field in the middle
Encouraged by the mass exchange series diagrams, a natural question to ask is what happens when we increase Zeeman strength in the middle further. A first step answer to the ask is we will meet a system with higher Hall conductance. For instance, after increasing Zeeman field strength to $\alpha=1.2$, we see from Fig. 30(a) that the Hall conductivity of the system becomes $-2e^{2}/h$ now. For the reason behind, we again look on the effective masses presented in Fig. 30(b), where a pair of gapless Dirac cones and one non-trivial gapped Dirac cone with mass sign reversal emerge, and essentially, from Eq. 38 and Eq. 42, they contribute synergistically to the Hall conductivity, i.e., $1/2+1/2+1=2$ units over $-e^{2}/h$. A careful trace over the effective mass evolution upon increasing $\alpha$ reveals that, at this time, $n=3$ band of $\chi=+$ closes and reopens the gap, during which an avoid crossing happens and forces it to exchange low energy mass with $n=1$ band of $\chi=+$, which leads to the result above.
@@ VI.2 Higher Chern Number Insulator
Still we firstly present an example shown in Fig. 31(a) as the Chern numbers of a uniformly magnetized TI film with total layer number $L_{z}=8$. The algorithm follows [143]. With increasing the uniform Zeeman strength $V$, the change of Chern numbers experiences three stages: For the relatively weak Zeeman field, the Chern number plateau increases step by step, from 00 to $8$, as revealed by red dots; For the Zeeman field with medium strength, the Chern number plateau drops from $8$ to $-8$ with $2$ as a step, illustrated by blue dots; Finally for the relatively strong Zeeman field, the Chern number plateau again increases from $-8$ to 00 one-by-one, shown by purple dots. Notice that under our parameter choice we have $m_{1}(\pi,0)\sim 2\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ and $m_{1}(\pi,\pi)\sim 4.3\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$.
@@
In our case, $|m_{n}(\Gamma)|<|m_{n}(X)|<|m_{n}(M)|$, and admittedly, all bulk bands $n\geq 2$ are trivial by which we mean $m_{n}(\Gamma/X/M)$ share the same sign, then focusing on one band and increasing $V$ from zero, we see that when $V$ just crosses $|m_{n}(\Gamma)|$, the band with $\chi m_{n}<0$ increases its Chern number from zero to one; continuing to increase $V$ so that it is bigger that $m_{n}(X)$, the corresponding Chern number reverses its sign from $1$ to $-1$; and finally when $V$ goes beyond the bandwidth $|m_{n}(M)|$, the band goes back to its trivial phase with zero Chern number. Notice that under our assumption $V>0$, the band $\bar{\chi}m_{n}>0$ is always trivial.
It is now clear that the sub-band mass-inversion at $\Gamma,\leavevmode\nobreak\ X$ and $M$ points are responsible for the change of Chern numbers, or equivalently the anomalous Hall plateaus with quantum units of conductance revealed in Fig. 31(a). As presented in Fig. 31(b), the masses $m_{n}(\bm{k})$ now share the property that $\max[m_{n}(\Gamma)]<\min[m_{l}(X)]$, $\max[m_{n}(X)]<\min[m_{l}(M)]$, as revealed by the green guidance lines. Then the Chern number change can be divided into three regions with increasing Zeeman field $V$ labelled in Fig. 31(a), i.e., the $\Gamma$-mass inverse region, the $X(Y)$-mass inverse region and the $M$-mass inverse region, without crossing among distinct regions. The physics happening in each region is exactly $L_{z}=8$ copies illustrated above with increasing $V$, i.e., the Chern number increases one-by-one in the $\Gamma$-region each time Zeeman field $V$ crosses some $|m_{n}(\Gamma)|$ and makes the band non-trivial, until it reaches its maximum $C_{\max}=L_{z}=8$, then decreases two-by-two in the $X$-region once $V$ gets bigger than some $|m_{n}(X)|$, where topological phase transition happens with both sides non-trivial, until bottom touching $C_{\min}=L_{z}-2L_{z}=-8$, and finally the Chern number goes back to zero step-by-step in the $M$-region as long as $V$ becomes bigger than some bandwidth $|m_{n}(M)|$ and makes corresponding band trivial again. The inverse process happens for an opposite Zeeman field, with Chern number reversing its sign.
@@ VI.3 Cooperation between middle and surfaces
Figure 33: (a) Hall conductivity of a metallic QAHE with its top and bottom surface states also gapped by magnetism, whose polarization direction is parallel to the field in the middle. (b) Momentum-dependent effective masses of Dirac fermions in Eq. (68). Due to the symmetric Zeeman configurations, masses are again divided into mirror classes by $\chi=\pm$. The masses for gapped surface states have been stressed in the same color. (c) The band structure of the system. Specifically, here the total layer number of TI film is $L_{z}=22$, the magnetic layer numbers at top, middle and bottom are $2$, $6$, $2$, with mean Zeeman strengths chosen to be $V^{\text{top}}=0.05\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, $V^{\text{mid}}=\alpha t_{\perp}$ with $\alpha=0.9$, and $V^{\text{bot}}=0.05\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, respectively.
@@
Now let us switch down $V^{\text{top}}$, which makes $V_{S}^{\text{top}}=-V_{A}^{\text{top}}>0$, accordingly we have $N^{+}=1,\leavevmode\nobreak\ C^{-}=1$, which corresponds to a system with Hall conductivity $3e^{2}/2h$. Further we re-add $V^{\text{top}}=-V^{\text{bottom}}<0$, which leads to $V_{S}^{\text{top}}=0,\leavevmode\nobreak\ V_{A}^{\text{top}}>0$, and we see $C^{+}=1,\leavevmode\nobreak\ C^{-}=0$, which makes the system a Chern insulator again with unit Chern number. Next we reverse $V^{\text{bottom}}$ to minus, and $V_{S}^{\text{top}}<0,\leavevmode\nobreak\ V_{A}^{\text{top}}=0$, which makes the system trivial with $C^{+}=C^{-}=0$. Finally, we switch down again $V^{\text{top}}$, and now $V_{S}^{\text{top}}=-V_{A}^{\text{top}}<0$, accordingly we have $N^{+}=1,\leavevmode\nobreak\ C^{-}=0$, which leaves half quantization of Hall conductivity in the system.
Totally, we see that there exist five more additional topologically distinct phases upon tuning surface magnetism of metallic QAHE, with Hall conductivities quantized into $2,3/2,1,1/2$ and 00 over quantum units, respectively. The topological properties of these additional phases can be easily verified by calculating their Hall conductivities, or reading from their effective mass pictures. The signs of Hall conductivities are inverted once we overturn magnetism at both surfaces and in the middle.
@@ VII Discussion and conclusion
Half quantum mirror Hall effect [Uncaptioned image] [Uncaptioned image] $N^{\text{mirror}}=1-(-1)=2$
Half quantized anomalous Hall effect [Uncaptioned image] [Uncaptioned image] $C=0$, $N=1$
Metallic quantized anomalous Hall effect [Uncaptioned image] [Uncaptioned image] $N=1+1=2$
@@
The Hall conductivity associated with the gapless and gapped Dirac fermions in the TI film are ±$\pm e^{2}/2h$ and 00, respectively. This results in a half-quantized topological phase, serving as a metallic partner to the insulating quantum spin Hall effect, namely, the half quantum mirror Hall effect in TI film itself with a mirror symmetry. The pairing feature of the gapless Dirac fermions in half quantum mirror Hall effect is summarized in Table 3. It is noteworthy that their existence here is not a result from the Nielsen-Ninomiya theorem, since they are two separable fermions in whole Brillouin zone; rather, it is the mirror symmetry along the opened direction of the TI film that requires the doubling — symmetric and antisymmetric.
@@
The formalism introduced here, involving the transformation of a confined spatially $(n+1)$D Dirac Hamiltonian into $n$D Dirac fermions through the construction of a local unitary matrix using solutions from a decomposed 1D Hamiltonian along the confined direction, can be generalized to arbitrary dimensions, with the aid of Clifford algebra. In particular, initiating from a 4D space modified Dirac equation, a unitary transformation yields a pair of gapless Dirac fermions effectively in 3D space. This extension holds the potential to enhance our comprehension of the chiral anomaly in the system[56, 57]. What is more, given that the high-energy components of the two Dirac fermions explicitly break the chiral symmetry, they are not obliged to be paired by violating conditions stipulated by the Nielsen-Ninomiya theorem. As a result, we can anticipate that when the constrained 4D Hamiltonian becomes ‘semi-magnetic’, a single gapless Dirac fermion will be observed, similar to that in half QAHE.
@@
In the presence of a uniform Zeeman field, the mass of each Dirac fermion in TI film is directly modified by a Zeeman field. By tuning the strength of magnetism, sub-band inversion happens step-by-step for each Dirac fermion, whose Chern character changes correspondingly. Summing those mass-modified Dirac fermions together gives a Chern insulator that carries jumping Hall conductance among integers in $[-L_{z},L_{z}]$ over the quantum unit $e^{2}/h$, with $L_{z}$ the total layer number.
@@ Acknowledgements.
Now considering finite size along $z$ direction with top and bottom surfaces located at ±$\pm\dfrac{L}{2}$, respectively, one would have boundary condition
@@
$\dfrac{m_{1}\xi_{2}}{m_{2}\xi_{1}}=\dfrac{\tan\xi_{2}L/2}{\tan\xi_{1}L/2}\vspace{5pt}$ (158a)
@@
$E_{+}=m_{0}(\bm{k})-t_{\perp}\frac{\xi_{1}^{2}g^{+}(\xi_{1})-\xi_{2}^{2}g^{+}(\xi_{2})}{g^{+}(\xi_{1})-g^{+}(\xi_{2})},\leavevmode\nobreak\ g^{+}(\xi)=\frac{\tan(\xi L/2)}{\xi},$ (159a)
$E_{-}=m_{0}(\bm{k})-t_{\perp}\frac{\xi_{1}^{2}g^{-}(\xi_{1})-\xi_{2}^{2}g^{-}(\xi_{2})}{g^{-}(\xi_{1})-g^{-}(\xi_{2})},\leavevmode\nobreak\ g^{-}(\xi)=\frac{1}{\tan(\xi L/2)\xi}.$ (159b)
In common sense, it is time taking $E_{\pm}$ into expressions of $\xi$s, together with the coefficients equations again and solve them. However, that not only is tricky but lacks of physical insight, and we shall change our perspective.
@@
$\begin{cases}f_{+}(z)=\dfrac{\cos(\xi_{1}z)}{\cos(\xi_{1}L/2)}-\dfrac{\cos(\xi_{2}z)}{\cos(\xi_{2}L/2)}\vspace{5pt}\\ f_{-}(z)=\dfrac{\sin(\xi_{1}z)}{\sin(\xi_{1}L/2)}-\dfrac{\sin(\xi_{2}z)}{\sin(\xi_{2}L/2)}\end{cases},$ (160)
@@
$\phi^{+}_{1}=C^{+}_{1}\begin{pmatrix}-is\lambda_{\perp}f_{+}\vspace{2pt}\\ t_{\perp}\eta_{1}f_{-}\end{pmatrix},\ E=E_{+},$ (163a)
@@
$\phi^{+}_{2}=C^{+}_{2}\begin{pmatrix}t_{\perp}\eta_{1}f_{-}\\ is\lambda_{\perp}f_{+}\end{pmatrix},\leavevmode\nobreak\ E=-E_{+},$ (165a)
$\phi^{-}_{2}=C^{-}_{2}\begin{pmatrix}t_{\perp}\eta_{2}f_{+}\\ -is\lambda_{\perp}f_{-}\end{pmatrix},\leavevmode\nobreak\ E=-E_{-},$ (165b)
@@
Solution above seems to give four solutions, mathematical restriction, however, tells that equations from different lines for the same set of coefficients must stand simultaneously, i.e., (163a)$\Leftrightarrow$(165b) and (163b)$\Leftrightarrow$(165a), which gives us two relations as
@@
$\begin{cases}m=E_{+}=m_{0}(\bm{k})-t_{\perp}\dfrac{\xi_{1}^{2}g(\xi_{1})-\xi_{2}^{2}g(\xi_{2})}{g(\xi_{1})-g(\xi_{2})},\\ g(\xi)=\dfrac{\tan(\xi L/2)}{\xi},\\ \eta=\dfrac{\xi_{1}^{2}-\xi_{2}^{2}}{\xi_{1}\cot(\xi_{1}L/2)-\xi_{2}\cot(\xi_{2}L/2)},\\ C=C^{+}_{1}=C^{+}_{2},\end{cases}$ (170)
@@
$\Phi_{1}$ $=\begin{pmatrix}\varphi(+)\\ 0\end{pmatrix}=C\begin{pmatrix}-i\lambda_{\perp}f_{+}\\ t_{\perp}\eta f_{-}\\ 0\\ 0\end{pmatrix},\leavevmode\nobreak\ \Phi_{2}=\begin{pmatrix}0\\ \chi(-)\end{pmatrix}=C\begin{pmatrix}0\\ 0\\ t_{\perp}\eta f_{-}\\ -i\lambda_{\perp}f_{+}\end{pmatrix},$ (171)
$\Phi_{3}$ $=\begin{pmatrix}\chi(+)\\ 0\end{pmatrix}=C\begin{pmatrix}t_{\perp}\eta f_{-}\\ i\lambda_{\perp}f_{+}\\ 0\\ 0\end{pmatrix},\leavevmode\nobreak\ \Phi_{4}=\begin{pmatrix}0\\ \varphi(-)\end{pmatrix}=C\begin{pmatrix}0\\ 0\\ i\lambda_{\perp}f_{+}\\ t_{\perp}\eta f_{-}\end{pmatrix},$
with energy $(m,-m,-m,m)$, respectively. Notice that $\Phi_{3,4}$ are chiral partners of $\Phi_{1,2}$ by $-i\tau_{y}$, respectively. To obtain $m$, a set of closed equations need to be solved
@@
$\xi_{\alpha}$ $=\sqrt{-\frac{F}{D}+(-1)^{\alpha-1}\frac{\sqrt{R}}{D}},\leavevmode\nobreak\ \alpha=1,2,$ (172b)

### v3

@@ VI.1.3 Stronger field in the middle
Encouraged by the mass exchange series diagrams, a natural question to ask is what happens when we increase Zeeman strength in the middle further. A first step answer to the ask is we will meet a system with higher Hall conductance. For instance, after increasing Zeeman field strength to $\alpha=1.2$, we see from Fig. 30(a) that the Hall conductivity of the system becomes $2e^{2}/h$ now. For the reason behind, we again look on the effective masses presented in Fig. 30(b), where a pair of gapless Dirac cones and one non-trivial gapped Dirac cone with mass sign reversal emerge, and essentially, from Eq. 38 and Eq. 42, they contribute synergistically to the Hall conductivity, i.e., $1/2+1/2+1=2$ units over $e^{2}/h$. A careful trace over the effective mass evolution upon increasing $\alpha$ reveals that, at this time, $n=3$ band of $\chi=+$ closes and reopens the gap, during which an avoid crossing happens and forces it to exchange low energy mass with $n=1$ band of $\chi=+$, which leads to the result above.
@@ VI.2 Higher Chern Number Insulator
Still we firstly present an example shown in Fig. 31(a) as the Chern numbers of a uniformly magnetized TI film with total layer number $L_{z}=8$. The algorithm follows [143]. With increasing the uniform Zeeman strength $V$, the change of Chern numbers experiences three stages: For the relatively weak Zeeman field, the Chern number plateau increases step by step, from 0 to $8$, as revealed by red dots; For the Zeeman field with medium strength, the Chern number plateau drops from $8$ to 8-8 with $2$ as a step, illustrated by blue dots; Finally for the relatively strong Zeeman field, the Chern number plateau again increases from 8-8 to 0 one-by-one, shown by purple dots. Notice that under our parameter choice we have $m_{1}(\pi,0)\sim 2\penalty 10000\ $\mathrm{e}\mathrm{V}$\$ and $m_{1}(\pi,\pi)\sim 4.3\penalty 10000\ $\mathrm{e}\mathrm{V}$\$.
@@
In our case, $|m_{n}(\Gamma)|<|m_{n}(X)|<|m_{n}(M)|$, and admittedly, all bulk bands $n\geq 2$ are trivial by which we mean $m_{n}(\Gamma/X/M)$ share the same sign, then focusing on one band and increasing $V$ from zero, we see that when $V$ just crosses $|m_{n}(\Gamma)|$, the band with $\chi m_{n}<0$ increases its Chern number from zero to one; continuing to increase $V$ so that it is bigger that $m_{n}(X)$, the corresponding Chern number reverses its sign from $1$ to 1-1; and finally when $V$ goes beyond the bandwidth $|m_{n}(M)|$, the band goes back to its trivial phase with zero Chern number. Notice that under our assumption $V>0$, the band $\bar{\chi}m_{n}>0$ is always trivial.
It is now clear that the sub-band mass-inversion at $\Gamma,\penalty 10000\ X$ and $M$ points are responsible for the change of Chern numbers, or equivalently the anomalous Hall plateaus with quantum units of conductance revealed in Fig. 31(a). As presented in Fig. 31(b), the masses $m_{n}(\bm{k})$ now share the property that $\max[m_{n}(\Gamma)]<\min[m_{l}(X)]$, $\max[m_{n}(X)]<\min[m_{l}(M)]$, as revealed by the green guidance lines. Then the Chern number change can be divided into three regions with increasing Zeeman field $V$ labelled in Fig. 31(a), i.e., the $\Gamma$-mass inverse region, the $X(Y)$-mass inverse region and the $M$-mass inverse region, without crossing among distinct regions. The physics happening in each region is exactly $L_{z}=8$ copies illustrated above with increasing $V$, i.e., the Chern number increases one-by-one in the $\Gamma$-region each time Zeeman field $V$ crosses some $|m_{n}(\Gamma)|$ and makes the band non-trivial, until it reaches its maximum $C_{\max}=L_{z}=8$, then decreases two-by-two in the $X$-region once $V$ gets bigger than some $|m_{n}(X)|$, where topological phase transition happens with both sides non-trivial, until bottom touching $C_{\min}=L_{z}-2L_{z}=-8$, and finally the Chern number goes back to zero step-by-step in the $M$-region as long as $V$ becomes bigger than some bandwidth $|m_{n}(M)|$ and makes corresponding band trivial again. The inverse process happens for an opposite Zeeman field, with Chern number reversing its sign.
@@ VI.3 Cooperation between middle and surfaces
Figure 33: (a) Hall conductivity of a metallic QAHE with its top and bottom surface states also gapped by magnetism, whose polarization direction is parallel to the field in the middle. (b) Momentum-dependent effective masses of Dirac fermions in Eq. (68). Due to the symmetric Zeeman configurations, masses are again divided into mirror classes by $\chi=\pm$. The masses for gapped surface states have been stressed in the same color. (c) The band structure of the system. Specifically, here the total layer number of TI film is $L_{z}=22$, the magnetic layer numbers at top, middle and bottom are $2$, 66, 22, with mean Zeeman strengths chosen to be $V^{\text{top}}=0.05\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, $V^{\text{mid}}=\alpha t_{\perp}$ with $\alpha=0.9$, and $V^{\text{bot}}=0.05\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, respectively.
@@
Now let us switch down $V^{\text{top}}$, which makes $V_{S}^{\text{top}}=-V_{A}^{\text{top}}>0$, accordingly we have $N^{+}=1,\penalty 10000\ C^{-}=1$, which corresponds to a system with Hall conductivity $3e^{2}/2h$. Further we re-add $V^{\text{top}}=-V^{\text{bottom}}<0$, which leads to $V_{S}^{\text{top}}=0,\penalty 10000\ V_{A}^{\text{top}}>0$, and we see $C^{+}=1,\penalty 10000\ C^{-}=0$, which makes the system a Chern insulator again with unit Chern number. Next we reverse $V^{\text{bottom}}$ to minus, and $V_{S}^{\text{top}}<0,\penalty 10000\ V_{A}^{\text{top}}=0$, which makes the system trivial with $C^{+}=C^{-}=0$. Finally, we switch down again $V^{\text{top}}$, and now $V_{S}^{\text{top}}=-V_{A}^{\text{top}}<0$, accordingly we have $N^{+}=1,\penalty 10000\ C^{-}=0$, which leaves half quantization of Hall conductivity in the system.
Totally, we see that there exist five more additional topologically distinct phases upon tuning surface magnetism of metallic QAHE, with Hall conductivities quantized into $2,3/2,1,1/2$ and 0 over quantum units, respectively. The topological properties of these additional phases can be easily verified by calculating their Hall conductivities, or reading from their effective mass pictures. The signs of Hall conductivities are inverted once we overturn magnetism at both surfaces and in the middle.
@@ VII Discussion and conclusion
Half quantum
mirror Hall effect
[Uncaptioned image] [Uncaptioned image] $N^{\text{mirror}}=1-(-1)=2$
Half quantized
anomalous Hall effect
[Uncaptioned image] [Uncaptioned image] $C=0$, $N=1$
Metallic quantized
anomalous Hall effect
[Uncaptioned image] [Uncaptioned image] $N=1+1=2$
@@
The Hall conductivity associated with the gapless and gapped Dirac fermions in the TI film are $\pm e^{2}/2h$ and 0, respectively. This results in a half-quantized topological phase, serving as a metallic partner to the insulating quantum spin Hall effect, namely, the half quantum mirror Hall effect in TI film itself with a mirror symmetry. The pairing feature of the gapless Dirac fermions in half quantum mirror Hall effect is summarized in Table 3. It is noteworthy that their existence here is not a result from the Nielsen-Ninomiya theorem, since they are two separable fermions in whole Brillouin zone; rather, it is the mirror symmetry along the opened direction of the TI film that requires the doubling — symmetric and antisymmetric.
@@
The formalism introduced here, involving the transformation of a confined spatially (n+1)(n+1)D Dirac Hamiltonian into nnD Dirac fermions through the construction of a local unitary matrix using solutions from a decomposed 1D Hamiltonian along the confined direction, can be generalized to arbitrary dimensions, with the aid of Clifford algebra. In particular, initiating from a 4D space modified Dirac equation, a unitary transformation yields a pair of gapless Dirac fermions effectively in 3D space. This extension holds the potential to enhance our comprehension of the chiral anomaly in the system[56, 57]. What is more, given that the high-energy components of the two Dirac fermions explicitly break the chiral symmetry, they are not obliged to be paired by violating conditions stipulated by the Nielsen-Ninomiya theorem. As a result, we can anticipate that when the constrained 4D Hamiltonian becomes ‘semi-magnetic’, a single gapless Dirac fermion will be observed, similar to that in half QAHE.
@@
In the presence of a uniform Zeeman field, the mass of each Dirac fermion in TI film is directly modified by a Zeeman field. By tuning the strength of magnetism, sub-band inversion happens step-by-step for each Dirac fermion, whose Chern character changes correspondingly. Summing those mass-modified Dirac fermions together gives a Chern insulator that carries jumping Hall conductance among integers in $L_{z},L_{z}]$ over the quantum unit $e^{2}/h$, with $L_{z}$ the total layer number.
@@ Acknowledgements.
Now considering finite size along $z$ direction with top and bottom surfaces located at $\pm\dfrac{L}{2}$, respectively, one would have boundary condition
@@
$\dfrac{m_{1}\xi_{2}}{m_{2}\xi_{1}}=\dfrac{\tan\xi_{2}L/2}{\tan\xi_{1}L/2}\vskip 5.0pt$ (158a)
@@
$E_{+}=m_{0}(\bm{k})-t_{\perp}\frac{\xi_{1}^{2}g^{+}(\xi_{1})-\xi_{2}^{2}g^{+}(\xi_{2})}{g^{+}(\xi_{1})-g^{+}(\xi_{2})},\penalty 10000\ g^{+}(\xi)=\frac{\tan(\xi L/2)}{\xi},$ (159a)
$E_{-}=m_{0}(\bm{k})-t_{\perp}\frac{\xi_{1}^{2}g^{-}(\xi_{1})-\xi_{2}^{2}g^{-}(\xi_{2})}{g^{-}(\xi_{1})-g^{-}(\xi_{2})},\penalty 10000\ g^{-}(\xi)=\frac{1}{\tan(\xi L/2)\xi}.$ (159b)
[…]
@@
$\begin{cases}f_{+}(z)=\dfrac{\cos(\xi_{1}z)}{\cos(\xi_{1}L/2)}-\dfrac{\cos(\xi_{2}z)}{\cos(\xi_{2}L/2)}\vskip 5.0pt\\ f_{-}(z)=\dfrac{\sin(\xi_{1}z)}{\sin(\xi_{1}L/2)}-\dfrac{\sin(\xi_{2}z)}{\sin(\xi_{2}L/2)}\end{cases},$ (160)
@@
$\phi^{+}_{1}=C^{+}_{1}\begin{pmatrix}-is\lambda_{\perp}f_{+}\vskip 2.0pt\\ t_{\perp}\eta_{1}f_{-}\end{pmatrix},\ E=E_{+},$ (163a)
@@
$\phi^{+}_{2}=C^{+}_{2}\begin{pmatrix}t_{\perp}\eta_{1}f_{-}\\ is\lambda_{\perp}f_{+}\end{pmatrix},\penalty 10000\ E=-E_{+},$ (165a)
$\phi^{-}_{2}=C^{-}_{2}\begin{pmatrix}t_{\perp}\eta_{2}f_{+}\\ -is\lambda_{\perp}f_{-}\end{pmatrix},\penalty 10000\ E=-E_{-},$ (165b)
@@
[…]
@@
[…]
@@
$\Phi_{1}$ $=\begin{pmatrix}\varphi(+)\\ 0\end{pmatrix}=C\begin{pmatrix}-i\lambda_{\perp}f_{+}\\ t_{\perp}\eta f_{-}\\ 0\\ 0\end{pmatrix},\penalty 10000\ \Phi_{2}=\begin{pmatrix}0\\ \chi(-)\end{pmatrix}=C\begin{pmatrix}0\\ 0\\ t_{\perp}\eta f_{-}\\ -i\lambda_{\perp}f_{+}\end{pmatrix}$, (171)
$\Phi_{3}$ $=\begin{pmatrix}\chi(+)\\ 0\end{pmatrix}=C\begin{pmatrix}t_{\perp}\eta f_{-}\\ i\lambda_{\perp}f_{+}\\ 0\\ 0\end{pmatrix},\penalty 10000\ \Phi_{4}=\begin{pmatrix}0\\ \varphi(-)\end{pmatrix}=C\begin{pmatrix}0\\ 0\\ i\lambda_{\perp}f_{+}\\ t_{\perp}\eta f_{-}\end{pmatrix}$,
with energy (m,m,m,m)(m,-m,-m,m), respectively. Notice that $\Phi_{3,4}$ are chiral partners of $\Phi_{1,2}$ by $i\tau_{y}$, respectively. To obtain $m$, a set of closed equations need to be solved
@@
$\xi_{\alpha}$ $=\sqrt{-\frac{F}{D}+(-1)^{\alpha-1}\frac{\sqrt{R}}{D}},\penalty 10000\ \alpha=1,2,$ (172b)